# Mitigating Hallucination of Large Vision-Language Models via Dynamic Logits Calibration

### v1

@@ 5.3 Ablation Studies
We primarily validate the effectiveness of the adaptive factor in Dynamic Guidance Strength $\lambda_{t}$. Furthermore, in the Appendix, we explore the influence of the Modulation Strength $\alpha$, the impact of different Window Sizes $N$ (Appendix A.5), and the integration of DLC with various decoding methods (Appendix A.6).
Effectiveness of the adaptive factor in $\lambda_{t}$. To demonstrate the effectiveness of the adaptive factor $(1-\bar{B}_{t})^{2}$ in dynamic guidance strength $\lambda_{t}$, as formulated in Eq. 8, we configure models with and without this factor to generate responses on the COCO [28] dataset. The correctness ($C$) and detailedness ($D$) of these responses are evaluated using GPT-4o. As illustrated in Figure 7, models including this factor achieve higher scores in both accuracy and details compared to their counterparts without the factor, thereby validating its efficacy.
@@ Appendix A Appendix
This appendix provides supplementary materials to support the main findings of our paper. We begin with an overview of Large Vision-Language Models (LVLMs) in Section A.1, offering additional context to help readers understand the foundation of our research. Section A.2 presents a detailed pseudo-code of our proposed method, providing an in-depth algorithmic understanding. Section A.3 elaborates on the implementation details, including parameter settings and computational workflows. For a comprehensive evaluation of our approach, Section A.4 provides further details on the benchmarks used in our experiments. Section A.5 analyzes the impact of two key hyperparameters on performance: the modulation strength $\alpha$ and the historical context window size $N$. Section A.6 demonstrates the integration of DLC with various decoding strategies, including beam search, greedy search, top-k sampling, and temperature scaling. Section A.7 discusses the inherent limitations of our method, while Section A.8 presents additional comparative results of our approach against other methods on the LLaVA-Bench benchmark. Finally, Section A.9 showcases the prompts used for GPT-4 and GPT-4o evaluations.
@@ A.1 More backgrounds about LVLMs
The development of LVLMs has evolved from BERT-based [14, 33, 38] cross-modal alignment to LLM-powered [1, 7, 12, 26, 30, 56] architectures. Building upon foundational text-image alignment techniques like CLIP [35] and BLIP [25], contemporary LVLMs integrate visual encoders with LLM decoders through distinct optimization strategies. For instance, LLaVA [30] employs lightweight adapters combined with instruction tuning, while InstructBLIP [12] and MiniGPT-4 [56] utilize Q-Former modules for visual token compression. Despite these innovations, significant challenges persist, particularly severe hallucination issues arising from misaligned multimodal representations. To systematically analyze mitigation strategies, we have examined with three representative architectures, including LLaVA, InstructBLIP, and MiniGPT-4, whose results demonstrate that various technical approaches commonly share critical vulnerabilities to the hallucination issue.
@@ A.2 Pseudo-code of Our Method
3:$y\leftarrow x$triangleright Initialize output sequence with input prompt
4:$\bar{B}\leftarrow 0$triangleright Initialize historical baseline
@@
We compute all visual alignment scores using the SigLIP model [51]. The historical baseline $\bar{B}_{t}$ is calculated over a sliding window of $N=8$ past context scores, following a 3-step warm-up period. For the top-k candidates $(k=50)$ at each step, the combined score $S_{t,i}^{\text{COMB}}$ is computed. The adaptive modulation employs a maximum intervention strength $\alpha=3$, and the standard sigmoid function $\sigma(\cdot)$ for logit adjustment. Our experiments were run on a single A100 GPU.
@@ A.4 Further Details on Benchmarks
(1) CHAIR [36] quantifies object hallucination in image captions by measuring the ratio of objects mentioned in captions but absent from ground-truth labels. It includes two variants: CHAIRI ($C_{I}$), calculating hallucinated objects versus all generated objects, and CHAIRS ($C_{S}$), assessing sentences with hallucinations versus total sentences. These two metrics are defined as follows::
@@
(2) GPT-4 Assisted Evaluations [54] function as a crucial complement to conventional assessment methodologies. While traditional metrics such as CHAIR effectively quantify hallucinations at the object-existence level, they demonstrate inadequacy in detecting more fine-grained hallucination types, including positional, relational, and attribute-based inaccuracies. To address this limitation, GPT-4 assisted evaluation employs detailed object-level descriptions from the Visual Genome (VG) [20] dataset as ground-truth reference, leveraging GPT-4’s sophisticated capabilities for precise hallucination detection and classification. The complete prompt template is provided in the Appendix A.9.
(3) GPT-4o assisted evaluations [^1] We utilize GPT-4o for assisted evaluation, as it offers enhanced capabilities over previous vision-capable models like GPT-4V, which is also being phased out by OpenAI. [47] further conduct a comprehensive analysis of hallucinations and text quality in open-ended generation tasks. Harnessing GPT-4o’s sophisticated capabilities, [47] established a numerical assessment protocol utilizing a 0-10 scale, which evaluates two essential dimensions: Accuracy-the extent to which responses align with image content without fabrication, and Detailedness-the degree of informational richness present in the responses.
(4) LLaVA-Bench-in-the-Wild [30] is designed to evaluate LVLMs’ proficiency in addressing complex tasks and their adaptability to novel domains. We perform targeted case studies on this dataset to demonstrate the efficacy of our proposed DLC.
@@ A.5 Sensitivity Analysis of Key Hyperparameters: $\alpha$ and $N$
The performance of our DLC method is influenced by two key hyperparameters: the modulation strength $\alpha$ and the context window size $N$. In this section, we conduct a comprehensive sensitivity analysis to understand how these parameters affect the model’s hallucination mitigation capability and generation quality. Experiments were conducted using a range of values for these parameters across three distinct LVLMs, with performance assessed by the CHAIR benchmark [36] (max new tokens: 512 for all ablation studies).
@@
A.6 DLC Performance with Various Decoding Strategies
@@
Our proposed DLC demonstrates compatibility with various mainstream decoding strategies. To comprehensively evaluate its effectiveness, in addition to its performance with common strategies such as nucleus sampling and greedy sampling, we conducted further experiments on the LLaVA-1.5 [30] model using the MSCOCO dataset [28] and the CHAIR benchmark [36]. These experiments explored other decoding methods and parameters, specifically: Top-p sampling (p=0.9), Top-k sampling (k=50), and Top-k sampling (k=50) with varying temperatures (t=1.5 and t=0.8). The results, presented in Figure 8, consistently indicate that regardless of the sampling method employed, DLC effectively reduces hallucinations and enhances overall model performance. This substantiates the robustness and effectiveness of our proposed DLC across diverse decoding strategies.
@@
A.7 Discussion of Methodological Limitations and Social impacts.
@@
Scope of Applicability: The DLC method is primarily optimized for open-ended, descriptive text generation scenarios, aiming to mitigate semantic drift that may occur during the generation of lengthy texts by dynamically calibrating the probabilities of candidate tokens. Consequently, its advantages may not be fully realized in contexts requiring minimalistic answers, such as "yes/no" judgments (e.g., as encountered in tasks like the POPE benchmark [27]). The core mechanism of DLC relies on the generated textual context for dynamic adjustments; for extremely short answers, such as single "yes" or "no" responses, the lack of sufficient contextual information for effective logits calibration might constrain its performance. To address this limitation, future work could explore transforming such concise answers (e.g., "yes" or "no") into complete declarative sentences. For instance, for object presence verification, sentences like "There is an <object> in the image" versus "There is not an <object> in the image" could be constructed. Subsequently, CLIP could be employed to compute the semantic similarity between these statements and the input image, thereby assisting the model in selecting a more accurate response. This direction warrants further exploration and validation.
Dependency on CLIP Model Performance: The efficacy of DLC is intrinsically linked to the performance of the employed CLIP model. The CLIP model is tasked with assessing the semantic similarity between images and text, which serves as the foundational input for DLC’s logits calibration process. Therefore, if the CLIP model itself exhibits deficiencies or biases in accurately capturing the semantic associations for specific image-text pairs-stemming, for example, from biases in its training data or an insufficient understanding of certain complex, fine-grained visual concepts-this can directly impact the precision and overall effectiveness of the DLC calibration. Theoretically, leveraging more advanced models with superior vision-text alignment capabilities is anticipated to further enhance the performance ceiling of DLC.
@@
A.8 Extended Comparative Results on LLaVA-Bench
In this section, we present additional comparative results of our DLC approach against other methods on the LLaVA-Bench [30]. These extended evaluations provide a more comprehensive understanding of our method’s performance across various question types and scenarios.
A.9 Prompts for GPT-4 and GPT-4o Evaluation

### v2

@@ 5.3 Ablation Studies
We primarily validate the effectiveness of the adaptive factor $\lambda_{t}$. The Appendix provides a large number of additional experiments, including comprehensive ablation studies on key components and hyperparameters. These experiments also demonstrate DLC’s robustness across different CLIP models, its broader applicability to VQA tasks, and its scalability to larger 13B architectures.
Effectiveness of the adaptive factor in $\lambda_{t}$. To demonstrate the effectiveness of the adaptive factor $(1-\bar{B}_{t})^{2}$ in dynamic guidance strength $\lambda_{t}$, as formulated in Eq. 8, we configure models with and without this factor to generate responses on the COCO (Lin et al., 2014) dataset. The correctness ($C$) and detailedness ($D$) of these responses are evaluated using GPT-4o. As illustrated in Figure 7, models including this factor achieve higher scores in both accuracy and details compared to their counterparts without the factor, thereby validating its efficacy.
@@ Appendix A Appendix
This appendix provides supplementary materials to support the main findings of our paper. We begin with an overview of Large Vision-Language Models (LVLMs) in Section A.1, offering additional context to help readers understand the foundation of our research. Section A.2 presents a detailed pseudo-code of our proposed method, providing an in-depth algorithmic understanding. Section A.3 elaborates on the implementation details, including parameter settings and computational workflows. For a comprehensive evaluation of our approach, Section A.4 provides further details on the benchmarks used in our experiments. Section A.5 analyzes the impact of two key hyperparameters on performance: the modulation strength $\alpha$ and the historical context window size $N$. Section A.6 analyzes the robustness of DLC across different CLIP model variants. Section A.7 presents the extension of DLC to Visual Question Answering tasks, demonstrating its applicability beyond open-ended generation. Section A.8 provides comprehensive ablation studies validating the importance of each component in our framework. Section A.9 demonstrates DLC’s effectiveness on larger 13B-parameter models, confirming its scalability beyond 7B architectures. Section A.10 demonstrates the integration of DLC with various decoding strategies, including beam search, greedy search, top-k sampling, and temperature scaling. Section A.11 discusses the inherent limitations of our method, while Section A.13 presents additional comparative results of our approach against other methods on the LLaVA-Bench benchmark. Finally, Section A.12 showcases the prompts used for GPT-4 and GPT-4o evaluations.
@@ A.1 More backgrounds about LVLMs
The development of LVLMs has evolved from BERT-based [Devlin et al., 2019, Lu et al., 2019, Tan and Bansal, 2019] cross-modal alignment to LLM-powered [Alayrac et al., 2022, Chen et al., 2023, Dai et al., 2023, Li et al., 2023b, Liu et al., 2023, Zhu et al., 2023] architectures. Building upon foundational text-image alignment techniques like CLIP [Radford et al., 2021] and BLIP [Li et al., 2022], contemporary LVLMs integrate visual encoders with LLM decoders through distinct optimization strategies. For instance, LLaVA [Liu et al., 2023] employs lightweight adapters combined with instruction tuning, while InstructBLIP [Dai et al., 2023] and MiniGPT-4 [Zhu et al., 2023] utilize Q-Former modules for visual token compression. Despite these innovations, significant challenges persist, particularly severe hallucination issues arising from misaligned multimodal representations. To systematically analyze mitigation strategies, we have examined with three representative architectures, including LLaVA, InstructBLIP, and MiniGPT-4, whose results demonstrate that various technical approaches commonly share critical vulnerabilities to the hallucination issue.
@@ A.2 Pseudo-code of Our Method
[…]
@@
[…]
@@ A.4 Further Details on Benchmarks
(1) CHAIR Rohrbach et al. [2018] quantifies object hallucination in image captions by measuring the ratio of objects mentioned in captions but absent from ground-truth labels. It includes two variants: CHAIRI ($C_{I}$), calculating hallucinated objects versus all generated objects, and CHAIRS ($C_{S}$), assessing sentences with hallucinations versus total sentences. These two metrics are defined as follows::
@@
(2) GPT-4 Assisted Evaluations Zhao et al. [2023] function as a crucial complement to conventional assessment methodologies. While traditional metrics such as CHAIR effectively quantify hallucinations at the object-existence level, they demonstrate inadequacy in detecting more fine-grained hallucination types, including positional, relational, and attribute-based inaccuracies. To address this limitation, GPT-4 assisted evaluation employs detailed object-level descriptions from the Visual Genome (VG) Krishna et al. [2017] dataset as ground-truth reference, leveraging GPT-4’s sophisticated capabilities for precise hallucination detection and classification. The complete prompt template is provided in the Appendix A.12.
(3) GPT-4o assisted evaluations [^1] We utilize GPT-4o for assisted evaluation, as it offers enhanced capabilities over previous vision-capable models like GPT-4V, which is also being phased out by OpenAI. Yin et al. [2024] further conduct a comprehensive analysis of hallucinations and text quality in open-ended generation tasks. Harnessing GPT-4o’s sophisticated capabilities, [Yin et al., 2024] established a numerical assessment protocol utilizing a 0-10 scale, which evaluates two essential dimensions: Accuracy-the extent to which responses align with image content without fabrication, and Detailedness-the degree of informational richness present in the responses.
(4) LLaVA-Bench-in-the-Wild Liu et al. [2023] is designed to evaluate LVLMs’ proficiency in addressing complex tasks and their adaptability to novel domains. We perform targeted case studies on this dataset to demonstrate the efficacy of our proposed DLC.
@@ A.5 Sensitivity Analysis of Key Hyperparameters: $\alpha$ and $N$
The performance of our DLC method is influenced by two key hyperparameters: the modulation strength $\alpha$ and the context window size $N$. In this section, we conduct a comprehensive sensitivity analysis to understand how these parameters affect the model’s hallucination mitigation capability and generation quality. Experiments were conducted using a range of values for these parameters across three distinct LVLMs, with performance assessed by the CHAIR benchmark [Rohrbach et al., 2018] (max new tokens: 512 for all ablation studies).
@@
A.6 Robustness Analysis Across Different CLIP Models
To analyze the robustness of our DLC method and its dependence on the underlying CLIP model, we conducted comprehensive experiments using different CLIP backbone architectures. This analysis addresses concerns about the method’s reliance on a specific vision-language alignment model.
Table 6 presents the performance of DLC when integrated with various CLIP variants, including the original CLIP models, SigLIP, and the recent FG-CLIP model.
Table 6: Robustness Analysis Across Different CLIP Models. CHAIRS and CHAIRI scores (lower is better) for DLC when using different CLIP backbones. Best results are highlighted in bold.
CLIP Backbone LLaVA-1.5 InstructBLIP
[…]
CLIP-ViT-Large-Patch14 40.4 11.4 51.8 14.7
CLIP-ViT-Large-Patch14-336 39.7 10.6 51.2 14.4
SigLIP-SO400M-Patch14-384 39.4 10.9 51.0 14.3
FG-CLIP-Large [Xie et al., 2025] 37.8 10.4 49.8 13.9
The results demonstrate that while stronger CLIP models generally lead to better performance, DLC maintains consistent improvements across all tested architectures. This robustness indicates that our method is not brittlely tied to a specific CLIP variant, but rather benefits from the general principle of visual-textual alignment assessment. The superior performance with FG-CLIP, which specializes in fine-grained visual-textual alignment, particularly validates our approach’s potential for further enhancement as CLIP models continue to evolve.
A.7 Extension to Visual Question Answering Tasks
While DLC was initially designed for open-ended descriptive generation tasks, we have successfully extended it to Visual Question Answering (VQA) domains through simple modifications. By treating the user’s query content as already generated context, DLC demonstrates excellent performance on VQA tasks that require semantic-level reasoning.
We validated this extension on the popular POPE benchmark [Li et al., 2023c], which evaluates object existence in images through yes/no questions across three different setups: random, popular, and adversarial. The results are presented in Table 7.
The results demonstrate the effectiveness of DLC, validated by its successful extension from open-ended generation to structured VQA scenarios.
Table 7: Performance on POPE Benchmark. Results show DLC’s effectiveness when extended to VQA tasks. Accuracy (Acc) and F1 scores are reported for different setups. Best results are highlighted in bold.
Method Setup LLaVA-1.5 InstructBLIP
[…]
Sample Random 83.90 84.23 83.13 83.43
ICD 85.33 85.48 82.46 83.68
VCD 82.50 83.20 83.60 83.69
SID 87.56 87.10 84.36 84.43
DLC+SigLIP (Ours) 87.83 88.14 88.93 88.81
DLC+FG-CLIP (Ours) 88.87 88.88 89.53 89.32
Sample Popular 80.93 81.85 77.20 78.78
ICD 82.16 82.80 74.20 77.66
VCD 79.86 81.13 77.13 78.68
SID 84.16 84.11 77.10 78.98
DLC+SigLIP (Ours) 84.37 85.23 82.73 83.69
DLC+FG-CLIP (Ours) 85.33 85.84 82.53 83.42
Sample Adversarial 75.57 77.83 74.80 77.02
ICD 76.33 78.39 72.03 76.15
VCD 75.10 77.66 75.06 76.99
SID 79.60 80.43 76.03 78.07
DLC+SigLIP (Ours) 77.00 79.68 79.03 80.82
DLC+FG-CLIP (Ours) 78.70 80.68 79.57 81.10
A.8 Comprehensive Component Ablation Studies
To thoroughly validate the importance of each component in our proposed DLC framework, we conducted comprehensive ablation studies. We systematically evaluated the impact of removing individual components: Current Contextual Token Alignment (CCTA), Isolated Token Alignment (ITA), and Relative Visual Advantage (RVA).
The ablation experiments were conducted using GPT-4o assisted evaluation on the COCO dataset, with models generating detailed descriptions for 500 randomly selected images. Performance is assessed on two key dimensions: Correctness (accuracy of visual descriptions) and Detailedness (richness of information).
Table 8: Comprehensive Ablation Study Results. GPT-4o evaluation scores (0-10 scale, higher is better) demonstrating the importance of each DLC component. Best results are highlighted in bold.
Settings LLaVA-1.5 InstructBLIP
[…]
w/o CCTA 6.22 6.65 5.79 6.20
w/ CCTA 7.00 6.72 6.60 6.46
w/o ITA 6.22 6.61 5.84 6.30
w/ ITA 6.97 6.72 6.63 6.52
w/o RVA 6.35 6.51 5.83 6.25
w/ RVA 7.00 6.69 6.71 6.60
The results clearly demonstrate that each component contributes significantly to the overall performance of DLC.
A.9 Scalability to Larger Model Architectures
To validate the effectiveness of our method on larger-scale models, we conducted comprehensive experiments on 13B-parameter architectures, specifically evaluating DLC on LLaVA-1.5-13B and InstructBLIP-13B using both the CHAIR benchmark and GPT-4o assisted evaluation.
A.9.1 CHAIR Benchmark Results on 13B Models
Table 9 presents the CHAIR evaluation results for 13B models across different maximum token settings. The results consistently demonstrate that DLC provides substantial improvements over nucleus sampling baselines on both model architectures.
The results show remarkable improvements across both token length settings
Table 9: CHAIR Evaluation Results on 13B Models. Comparison of nucleus sampling baseline and DLC on larger model architectures. Lower CHAIR scores indicate better performance (fewer hallucinations). Best results are highlighted in bold.
Methods Max Token LLaVA-1.5-13B InstructBLIP-13B
[…]
Nucleus Sampling 64 23.20 8.16 27.80 10.22
DLC (Ours) 64 18.20 5.53 24.20 7.75
Nucleus Sampling 512 53.20 15.07 64.00 19.57
DLC (Ours) 512 33.00 8.80 55.60 15.37
A.9.2 GPT-4o Assisted Evaluation on 13B Models
To complement the CHAIR evaluation, we conducted GPT-4o assisted evaluation focusing on correctness and detailedness of generated descriptions. The results are presented in Table 10.
The GPT-4o evaluation demonstrates substantial improvements in both correctness and detailedness.
Table 10: GPT-4o Evaluation Results on 13B Models. Performance comparison on correctness and detailedness metrics. Best results are highlighted in bold.
Settings LLaVA-1.5-13B InstructBLIP-13B
[…]
Nucleus Sampling 5.25 6.05 4.26 5.06
DLC (Ours) 7.75 6.77 7.00 6.36
A.10 DLC Performance with Various Decoding Strategies
@@
Our proposed DLC demonstrates compatibility with various mainstream decoding strategies. To comprehensively evaluate its effectiveness, in addition to its performance with common strategies such as nucleus sampling and greedy sampling, we conducted further experiments on the LLaVA-1.5 [Liu et al., 2023] model using the MSCOCO dataset [Lin et al., 2014] and the CHAIR benchmark [Rohrbach et al., 2018]. These experiments explored other decoding methods and parameters, specifically: Top-p sampling (p=0.9), Top-k sampling (k=50), and Top-k sampling (k=50) with varying temperatures (t=1.5 and t=0.8). The results, presented in Figure 8, consistently indicate that regardless of the sampling method employed, DLC effectively reduces hallucinations and enhances overall model performance. This substantiates the robustness and effectiveness of our proposed DLC across diverse decoding strategies.
@@
A.11 Discussion of Methodological Limitations and Social impacts.
@@
Limited Applicability to Spatial and Coordinate-Level Tasks: While our extension to VQA tasks shows promising results on object existence verification (POPE benchmark [Li et al., 2023c]), DLC faces inherent limitations with tasks requiring precise spatial reasoning or coordinate-level grounding. Our method relies on CLIP’s global semantic assessment, which lacks sensitivity to fine-grained spatial details or numerical coordinates. This limitation affects tasks such as: (1) object counting and spatial relationship reasoning (e.g., "How many red cars are to the left of the blue building?"), (2) precise coordinate grounding tasks like object detection with bounding box generation, and (3) referring expression comprehension requiring exact localization.
Dependence on Visual Content Quality: DLC’s effectiveness is inherently tied to the quality and clarity of the input visual content. For images with poor quality, ambiguous visual information, the method’s ability to provide reliable calibration signals may be compromised. In such cases, the visual alignment scores from CLIP may not provide sufficiently discriminative information to guide token selection effectively, potentially limiting the method’s benefits.
@@
A.12 Prompts for GPT-4 and GPT-4o Evaluation
@@
A.13 Extended Comparative Results on LLaVA-Bench
In this section, we present additional comparative results of our DLC approach against other methods on the LLaVA-Bench [Liu et al., 2023]. These extended evaluations provide a more comprehensive understanding of our method’s performance across various question types and scenarios.